7.49. A pipelined processor has the following characteristics: - Loads 18% . Load stall (load penalty) 1 cycle Branches 2296 Probability a branch is taken 80% e Branch penalty on taken 3 cycles . RAW penalty 1 cycle Estimate the average cycles per instruction for this processor. RAW dependencies 20% of all instructions except branches
Answer #1

Solution:

When the branch is taken then the number of cycles = 0.22 * (0.8 * 3 + 0.2 * 1)

= 0.572 cycle

rest 78% will take 1 cycleso, 0.78

Total = 0.572 + 0.78 = 1.352 cycle

I hope this helps if you find any problem. Please comment below. Don't forget to give a thumbs up if you liked it. :)